Consider a bond with a coupon rate of 10% and annual coupons. The par value is $1,000, and the bond has 5 years to maturity. The yield to maturity is 9%. What is the value of the bond? (in dollars)

987

1008

1006

918

1087

1020

996

1039

971

956
Answer #1

Price of the bond is calculated by solving the below equation:

F Price = Sum of Discounted coupons +7 (1 + r) 1- (1+r)-n 1000 Price = Coupon x = *(1 +0.09)5 1-(1+0.09) -5 1000 Price = (0.

This can be verified from the below schedule:

Year CF Discount Factor Discounted CF
0 $             -   1/(1+0.09)^0= 1 1*0= $              -  
1 $    100.00 1/(1+0.09)^1= 0.917431193 0.91743119266055*100= $       91.74
2 $    100.00 1/(1+0.09)^2= 0.841679993 0.84167999326656*100= $       84.17
3 $    100.00 1/(1+0.09)^3= 0.77218348 0.772183480061064*100= $       77.22
4 $    100.00 1/(1+0.09)^4= 0.708425211 0.708425211065196*100= $       70.84
5 $1,100.00 1/(1+0.09)^5= 0.649931386 0.649931386298345*1100= $     714.92
Price= Sum of all Discounted CF $ 1,038.90